I really need help with this one .Please can you e
I really need help with this one .Please can you explain me with equations as well .
Answer #1

1)

magnetic field by a long current carrying wire is given as

B = (\mu/4\pi) (2I1/r)

magnetic field at A is given as

BA = (\mu/4\pi) (2I1/rA)

at A , rA = 2 m

BA = (10-7) (2 x 20 /2) = 2 x 10-6 T                inward to the page

magnetic field at B is given as

BB = (\mu/4\pi) (2I1/rB)

at B , rB = 12 m

BB = (10-7) (2 x 20 /12) = 0.33 x 10-6 T inward to the page

b)

FAC = (\mu/4\pi) (2I1I2 (AC)/rA) = (10-7) (2 x 20 x 50 x 15 / 2) = 1.5 x 10-3 N towards left

FBD = (\mu/4\pi) (2I1I2 (BD)/rB) = (10-7) (2 x 20 x 50 x 15 / 12) = 0.25 x 10-3 N    towards right

C)

Net force = F = FAC - FBD = 1.5 x 10-3 - 0.25 x 10-3 = 1.25 x 10-3    towards left

d)

dA = small area of rectangular strip = 15 dr

B = magnetic field through the strip = (\mu/4\pi) (2I1/r)

small magnetic flux is given as

d\phi = B dA = (\mu/4\pi) (2I1 15 dr/r)

d\phi = (\mu/4\pi) (2I1 15) (dr/r)

integrating both side

\phi = (\mu/4\pi) (2I1 15) (lnrB - lnrA) = (10-7) (2 x 20 x 15) (ln12 - ln2) = 1.075 x 10-4